# Multilinear Wavelet Compact T(1) Theorem

### v1

@@ Abstract.
A. W. Green’s research supported in part by NSF grant NSF-DMS-2202813.
@@ 1. Introduction
An $(m+1)$-linear form $\Lambda$ defined on the $(m+1)$-fold product of the Schwarz space ${\mathcal{S}}(\mathbb{R}^{d})$ is a singular integral form if its off-diagonal kernel satisfies the standard size and smoothness estimates (see Definition 3.1 below). A singular integral form $\Lambda$ is bounded on $L^{p_{0}}(\mathbb{R}^{d})\times\cdots\times L^{p_{m}}(\mathbb{R}^{d})$ for all
@@
where $\Lambda^{*,j}$ permutes the 00th and $j$th argument (see (1.3) below). When $m=1$, this is the celebrated $T(1)$ theorem of David and Journé [david84] which was extended to $m\geq 2$ by Grafakos and Torres [grafakos-torres] (see also the earlier works [christ1987polynomial, coifman-meyer]) and more recently developed in, for example, [DPLiMV2, DPLiMV1, lerner2019intuitive, li2019bilinear, li2014sharp].
Our goal here is to prove a $T(1)$ theorem for compactness of multilinear singular integral forms. The first distinction between the boundedness problem and compactness problem is that compactness is a property of operators, while boundedness in the reflexive range (1.1) can be equivalently stated in terms of forms. Accordingly for each $j=0,1,\ldots,m$, we associate to $\Lambda$ the $m$-linear adjoint operators $T^{*,j}$ and transposed forms $\Lambda^{*,j}$ by
@@
If $\Lambda$ is Calderón-Zygmund, then we say each $T^{*,j}$ is an $m$-linear Calderón-Zygmund operator (CZO). Furthermore, define $\mathsf{T}_{\Lambda}=\{T^{*,j}\}_{j=0}^{m}$.
@@
In the linear case, Villaroya in [Vil15] gave a complete characterization of compact Calderón-Zygmund operators on $L^{2}(\mathbb{R}^{d})$, which was further developed in [OV, PPV, SVW, Vil19]. Recently, Mitkovski and Stockdale in [MS23] gave a simplified formulation of the $T(1)$ theorem for compactness of Villaroya. More precisely, they showed that a CZO $T$ is compact if and only if $\mathfrak{b}_{0}$ and $\mathfrak{b}_{1}$ both belong to $\mathrm{CMO}(\mathbb{R}^{d})$ and a vanishing version of the weak boundedness property, called the weak compactness property, is satisfied. In contrast, a multilinear version of these compactness testing theorems remains unexplored. These $T(1)$ theorems for compactness are complemented by a compact Rubio De Francia theory of extrapolation developed by Hytönen and Lappas in [HL2, HL1]. Subsequently, in [COY], Cao, Olivo, and Yabuta extended the bilinear results of [HL2] to the multilinear setting and to the quasi-Banach range, in which case the target space can be $L^{r}$ with $r>\frac{1}{m}$. However, due to the difficulties of extrapolating to the upper endpoint in the multilinear setting [li2021endpoint, nieraeth2019quantitative], the results of [COY, HL2] do not consider the case where one (or more) input spaces is $L^{\infty}(\mathbb{R}^{d})$. We point out that our results below do yield compactness when an input space is $L^{\infty}(\mathbb{R}^{d})$.
@@
To state our main result in the bilinear setting (see Theorem 2 below for the full $m$-linear result), it remains only to introduce our bilinear version of the weak compactness property. Given $\varphi\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$, $w\in\mathbb{R}^{d}$, and $t>0$, define
@@
Let $\Lambda$ be a trilinear Calderón-Zygmund form and $\phi$ be a mother wavelet. Then, there exists $\psi^{1},\psi^{2}\in{\mathcal{S}}(\mathbb{R}^{d})$ such that the following are equivalent:
@@
    Each $T\in\mathsf{T}_{\Lambda}$, $1<p_{1},p_{2}\leq\infty$, and $\tfrac{1}{2}<r<\infty$ satisfying $\frac{1}{p_{1}}+\frac{1}{p_{2}}=\frac{1}{r}$, $T$ extends to a compact bilinear operator from $L^{p_{1}}(\mathbb{R}^{d})\times L^{p_{2}}(\mathbb{R}^{d})$ to $L^{r}(\mathbb{R}^{d})$ and from $L^{\infty}_{0}(\mathbb{R}^{d})\times L^{\infty}_{0}(\mathbb{R}^{d})$ to $\mathrm{CMO}(\mathbb{R}^{d})$.
@@
    For each $T\in\mathsf{T}_{\Lambda}$, $T(1,1)\in\mathrm{CMO}(\mathbb{R}^{d})$, and $\lim\limits_{z\to\infty}\mathsf{W}_{\Lambda}(z)=0$.
If the testing conditions in B. of Theorem 1 are satisfied, then we say $\Lambda$ is a compact Calderón-Zygmund form. Aside from having simple and precise wavelet testing conditions that ensure compactness of the adjoint operators, we also have a concrete description of compact Calderón-Zygmund forms. We refer to Theorem 2 and Definitions 2.6 and 5.1 below for a precise statement, but in words, we will show that a compact CZ form is a finite linear combination of compact wavelet forms and compact paraproduct forms. The strategy of representing CZOs as simpler model operators has been very successful in understanding more refined properties of CZOs, most famously in providing sharp quantitative weighted estimates. Dyadic representation theorems, finding their roots in the works of Figiel [figiel90]; Nazarov, Treil, and Volberg [NTV2]; and Petermichl [PetAJM] among others, have led to a powerful and comprehensive theory of quantitative $L^{p}$ estimates for singular integrals, culminating with Hytönen’s proof of the $A_{2}$ conjecture [hytonen12]. See [m2022, HL2022, li2019bilinear, martikainen2012representation] and references therein for more recent extensions. Wavelet representations, which may be traced back to David and Journé’s original proof of the $T(1)$ theorem, have recently been developed by Di Plinio and some of the authors in a series of papers [DGW2, DGW, DWW] primarily to extend the Lebesgue theory to Sobolev spaces, though some new Lebesgue space results have also been obtained as a byproduct.
Our main result is another example of a new result in the Lebesgue setting obtained through wavelet representations. This approach differs from the ones in [MS23] and [Vil15] as we do not use the machinery regarding localization, and neither do we require additional decay of the kernel, but rather we use a wavelet averaging procedure to obtain the representation formula in §3 and then use the Riesz-Kolmogorov criterion to obtain the precompactness of the image of the unit ball under the adjoint operators to the wavelet and paraproduct forms in §4 and §5.
@@ 2.1. Analysis of the parameter space
$|\mathbf{w}-w_{0}|_{2}=\sqrt{\sum_{i=1}^{m}|w_{i}-w_{0}|^{2}},\quad\mathbf{w}=(w_{1},\ldots,w_{m}),\quad w_{i}\in\mathbb{R}^{d}.$
For $\mathbf{z}=(\mathbf{w},s)\in\mathbb{Z}^{md}$ and $\zeta=(w_{0},t)\in\mathbb{Z}^{d}$, define
@@ 2.2. Wavelet classes and forms
$\Psi^{m,\delta}_{z}=\left\{\varphi\in\mathcal{C}^{\delta}(\mathbb{R}^{md}):\|({\mathsf{Sy}}_{\mathbf{z}})^{-1}\varphi\|_{\star,\delta}\leq 1\right\},\quad\mathbf{z}=(w,\ldots,w,t)\in Z^{md},$
@@
Furthermore, if $\phi_{j}\in{\mathcal{S}}^{0}$, then $\nu_{\zeta}\in C\Psi^{m,\eta;j}_{\zeta}$.
@@
Wavelet forms may be viewed as a generalization of the Calderón-Toeplitz operators considered in [nowak1993calderon, rochberg1990toeplitz]. More generally though, all cancellative CZ forms can be realized as wavelet forms [DGW, DWW]. Bounded wavelet forms are bounded in the following sense. In fact, they themselves are cancellative CZ forms so Proposition 2.8 below follows from any number of results [grafakos-torres, lerner2019intuitive, li2021endpoint, li2014sharp, nieraeth2019quantitative]. See also [DGW]*Proposition 5.1 for a direct proof of the sparse $(1,\ldots,1)$ bound for bounded wavelet forms. Let us introduce some bookkeeping to concisely describe the full range of Lebesgue space estimates for CZOs.
@@
$P=\left\{(p_{1},\ldots,p_{m}):1<p_{j}\leq\infty\right\}.$
@@
$Q=\left\{(\vec{p},r(\vec{p})):\vec{p}\in P,\ r(\vec{p})<\infty\right\}.$
@@
$\left\|T\right\|_{\infty,\mathrm{BMO}}=\sup_{(f_{1},\ldots,f_{m})\in\mathcal{B}^{\vec{p}}}\left\|T(f_{1},\ldots,f_{m})\right\|_{\mathrm{BMO}(\mathbb{R}^{d})},\quad\vec{p}=(\infty,\ldots,\infty).$
@@
For each $(\vec{p},r)\in Q$, there exists $C_{\vec{p},r}>0$ such that for any bounded wavelet form $U$ and any $T\in\mathsf{T}_{U}$,
@@
Furthermore, there exists $C_{\infty}>0$ such that
(2.8) $\left\|T\right\|_{\infty,\mathrm{BMO}}\leq C_{\infty}\sup_{z\in Z^{d}}\left|\varepsilon_{z}\right|.$
@@ 3. Wavelet representation of compact Calderón-Zygmund forms
$|K(x_{0},\ldots,x_{m})|\leq\dfrac{C_{K}}{(\sum_{j=1}^{m}|x_{0}-x_{j}|)^{md}},$
$\max_{j=0,\ldots,m}|\Delta^{j}_{h}K(x_{0},\ldots,x_{m})|\leq\dfrac{C_{K}|h|^{\delta}}{(\sum_{j=1}^{m}|x_{0}-x_{j}|)^{md+\delta}},$
were $\Delta_{h}^{j}$ denotes the difference operator in the $j$-th position. We say $\Lambda$ is an $m$-linear singular integral form if there exists a singular integral kernel $K$ such that for any $f_{0},\ldots,f_{m}\in{\mathcal{S}}$ with $\cap_{j=0}^{m}{\mathrm{supp}\,}f_{j}=\varnothing$ one has
@@
We need the following function spaces in order to define Calderón-Zygmund forms. For $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$, define the $\mathrm{BMO}(\mathbb{R}^{d})$ norm
$\left\|f\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}=\sup_{Q\,\mathrm{cube}}\frac{1}{|Q|}\int_{Q}\left|f(x)-f_{Q}\right|\,{\rm d}x,\quad f_{Q}=\frac{1}{|Q|}\int_{Q}f(y)\,{\rm d}y.$
Then, the functions of bounded mean oscillation (BMO) are those with finite $\mathrm{BMO}(\mathbb{R}^{d})$ norm. Let ${\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ be the space of all continuous functions $f$ on $\mathbb{R}^{d}$ for which $\lim_{|x|\to\infty}f(x)=0$. Then, the Banach space of functions with continuous mean oscillation, $\mathrm{CMO}(\mathbb{R}^{d})$, is defined to be the closure of ${\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ in the norm $\left\|\cdot\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}$. Furthermore, $\mathrm{BMO}(\mathbb{R}^{d})$ becomes a a Banach space upon identifying functions which differ by a constant.
@@
and furthermore, there exist $\mathfrak{b}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$ such that (1.2) holds. The rigorous definition of (1.2) is as follows. Let $\theta\in{\mathrm{C}}^{\infty}_{0}(\mathbb{R}^{d})$ with $\theta=1$ near the origin. Then, for $t>0$, set $\theta_{t}=\theta(\cdot t)$. We say (1.2) holds, and sometimes use the language $T^{*,j}(1,\ldots,1)=\mathfrak{b}_{j}$, if for all $\varphi\in{\mathcal{S}}^{0}(\mathbb{R}^{d})$,
@@
Furthermore, let $\phi$ be a mother wavelet and $\psi^{2},\psi^{4}$ be the Schwartz functions from Lemma 2.4. For $j=1,\ldots,m$ define

### v2

@@ Abstract.
A. W. Green’s research supported in part by NSF-DMS-2202813.
@@ 1. Introduction
A continuous $(m+1)$-linear form $\Lambda$ defined on the $(m+1)$-fold product of the Schwarz space ${\mathcal{S}}(\mathbb{R}^{d})$ is a singular integral form if its off-diagonal kernel satisfies the standard size and smoothness estimates (see Definition 3.1 below). A singular integral form $\Lambda$ is bounded on $L^{p_{0}}(\mathbb{R}^{d})\times\cdots\times L^{p_{m}}(\mathbb{R}^{d})$ for all
@@
where $\Lambda^{*,j}$ permutes the 00th and $j$th argument (see (1.3) below). When $m=1$, this is the celebrated $T(1)$ theorem of David and Journé [david84] which was extended to $m\geq 2$ by Grafakos and Torres [GT02] (see also the earlier works [christ1987polynomial, coifman-meyer]) and more recently developed in, for example, [DPLiMV2, DPLiMV1, LN19, li2019bilinear, li2014sharp].
Our goal here is to prove a $T(1)$ representation theorem for compactness of multilinear singular integral forms. The first distinction between the boundedness problem and compactness problem is that compactness is a property of operators, while boundedness in the reflexive range (1.1) can be equivalently stated in terms of forms. Accordingly for each $j=0,1,\ldots,m$, we associate to $\Lambda$ the $m$-linear adjoint operators $T^{*,j}$ and transposed forms $\Lambda^{*,j}$ by
@@
If $\Lambda$ is Calderón-Zygmund, then we say each $T^{*,j}$ is an $m$-linear Calderón-Zygmund operator (CZO). Often we use the shorthand $T=T^{*,0}$.
@@
In the linear case, Villaroya in [Vil15] gave a complete characterization of compact Calderón-Zygmund operators on $L^{2}(\mathbb{R}^{d})$, which was further developed in [OV, PPV, SVW, Vil19]. Recently, Mitkovski and Stockdale in [MS23] gave a simplified formulation of the $T(1)$ theorem for compactness of Villaroya. More precisely, they showed that a CZO $T$ is compact if and only if $\mathfrak{b}_{0}$ and $\mathfrak{b}_{1}$ both belong to $\mathrm{CMO}(\mathbb{R}^{d})$ and a vanishing version of the weak boundedness property, called the weak compactness property, is satisfied. Even more recently, Bényi, Li, Oh, and Torres [BLOT24b] have characterized compact CZOs in terms of $L^{2}$ and $L^{\infty}$ testing conditions, in the spirit of Stein’s formulation of the $T(1)$-theorem [stein-book]. These $T(1)$ theorems for compactness are complemented by a compact Rubio De Francia theory of extrapolation due to Hytönen and Lappas in [HL2, HL1], and subsequently developed by others in [COY, CST]. Particularly, in [COY], Cao, Olivo, and Yabuta extended the bilinear results of [HL2] to the multilinear setting and to the quasi-Banach range, in which case the target space can be $L^{r}$ with $r>\frac{1}{m}$. However, due to the difficulties of extrapolating to the upper endpoint in the multilinear setting [li2021endpoint, nieraeth2019quantitative], the results of [COY, HL2] do not consider the case where one (or more) input spaces is $L^{\infty}(\mathbb{R}^{d})$. We point out that our results below do yield compactness when an input space is $L^{\infty}(\mathbb{R}^{d})$.
Our main result is Theorem 2 below which builds on the literature discussed above in the following two noteworthy manners. First, it provides a $T(1)$-type characterization of compact multilinear CZOs. Though extrapolation of compactness had been developed in the multilinear setting as mentioned above, a $T(1)$-type characterization in the vein of Villaroya [Vil15] or the simplified formulation of Mitkovski and Stockdale [MS23] was heretofore unknown. Our characterization is indeed in the simplified form of [MS23] (not relying on compact CZ kernel estimates) while at the same time provides a stronger result than both [BLOT24b] and [MS23] when $m=1$.
Second, we provide a representation formula for all compact CZOs. This is one answer to the search for examples of compact CZOs. Frankly, from Theorem 2, we now know all the examples of $m$-linear compact CZOs. This contribution is new even in the linear case. In words, we can describe our wavelet representation formula in terms of two building blocks whose adjoint operators are well-known examples of compact CZOs. The first class of buiding blocks consists of what we call compact wavelet forms which are almost-orthogonal perturbations of compact approximations of the wavelet resolution of the identity. The second class of building blocks consists of paraproduct forms with $\mathrm{CMO}(\mathbb{R}^{d})$ symbol. The wavelet representation theorem we will prove says that all compact CZOs are finite linear combinations of these two building blocks. We refer to Theorem 2 and Definitions 2.6 and 5.1 below for precise statements.
The strategy of representing CZOs as simpler model operators has been very successful in understanding more refined properties of CZOs, most famously in providing sharp quantitative weighted estimates. Dyadic representation theorems, finding their roots in the works of Figiel [figiel90]; Nazarov, Treil, and Volberg [NTV2]; and Petermichl [PetAJM] among others, have led to a powerful and comprehensive theory of quantitative $L^{p}$ estimates for singular integrals, culminating with Hytönen’s proof of the $A_{2}$ conjecture [hytonen12]. See [m2022, HL2022, li2019bilinear, martikainen2012representation] and references therein for more recent extensions. Smooth wavelet representations, which may be traced back to David and Journé’s original proof of the $T(1)$ theorem, have recently been developed by Di Plinio and some of the authors in a series of papers [DGW2, DGW, DWW] in order to study various Sobolev space estimates for CZOs.
@@
For simplicity, here in the introduction we state only the bilinear testing characterization, without the representation. It remains only to introduce the “weak compactness property” in the bilinear setting. For the function $\varphi\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$, $w\in\mathbb{R}^{d}$, and $t>0$, define
@@
Let $T$ be a bilinear CZO, $\Lambda$ the associated trilinear form, and $\mathfrak{b}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$ defined by (1.2). Then, the following are equivalent.
@@
    For each $1<p_{1},p_{2}\leq\infty$, and $\tfrac{1}{2}<r<\infty$ satisfying $\frac{1}{p_{1}}+\frac{1}{p_{2}}=\frac{1}{r}$, $T$ extends to a compact bilinear operator from $L^{p_{1}}(\mathbb{R}^{d})\times L^{p_{2}}(\mathbb{R}^{d})$ to $L^{r}(\mathbb{R}^{d})$ and from $L^{\infty}(\mathbb{R}^{d})\times L^{\infty}(\mathbb{R}^{d})$ to $\mathrm{CMO}(\mathbb{R}^{d})$.
@@
    $\lim\limits_{z\to\infty}\mathsf{W}_{\Lambda}(z)=0$ and for each $j=0,1,2$, $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$.
A broad outline of the proof is to use a wavelet averaging procedure to obtain the representation formula in §3 and then use the Riesz-Kolmogorov criterion to obtain the precompactness of the image of the unit ball under the adjoint operators to the wavelet and paraproduct forms in §4 and §5. These results are combined in §6 to state our main result, Theorem 2 including the representation theorem and a simplified statement in the linear case, Theorem 3. There we also outline a sample application.
@@ 2.1. Analysis of the parameter space
(2.1) $|\mathbf{w}-w_{0}|_{2}=\sqrt{\sum_{i=1}^{m}|w_{i}-w_{0}|^{2}},\quad\mathbf{w}=(w_{1},\ldots,w_{m}),\quad w_{i}\in\mathbb{R}^{d}.$
For $\mathbf{z}=(\mathbf{w},s)\in\mathbb{Z}^{md}$ and $\zeta=(w_{0},t)\in\mathbb{Z}^{d}$, define for any $0<\delta\leq 1$,
@@ 2.2. Wavelet classes and forms
$\Psi^{m,\delta}_{z}=\left\{\varphi\in{\mathrm{C}}^{\delta}(\mathbb{R}^{md}):\|({\mathsf{Sy}}_{\mathbf{z}})^{-1}\varphi\|_{\star,\delta}\leq 1\right\},\quad\mathbf{z}=(w,\ldots,w,t)\in Z^{md},$
@@
Furthermore, if $\varphi_{j}\in{\mathcal{S}}^{0}(\mathbb{R}^{d})$, then $\nu_{\zeta}\in C\Psi^{m,\eta;j}_{\zeta}$.
@@
Wavelet forms may be viewed as a generalization of the Calderón-Toeplitz operators considered in [nowak1993calderon, rochberg1990toeplitz]. More generally though, all cancellative CZ forms (see Definition 3.2 below) can be realized as wavelet forms [DGW, DWW]. Let us proceed to describe in what sense bounded wavelet forms are indeed bounded. They themselves are cancellative CZ forms so Proposition 2.8 below follows from any number of results [GT02, LN19, li2021endpoint, li2014sharp, nieraeth2019quantitative]. See also [DGW]*Proposition 5.1 for a direct proof of the sparse $(1,\ldots,1)$ bound for bounded wavelet forms. Let us introduce some bookkeeping to concisely describe the full range of Lebesgue space estimates for CZOs.
@@
$P=\left\{(p_{1},\ldots,p_{m}):1<p_{j}\leq\infty\right\},\quad{P_{\circ}=\left\{(p_{1},\ldots,p_{m}):1<p_{j}<\infty\right\}}.$
@@
$Q=\left\{(\vec{p},r(\vec{p})):\vec{p}\in P,\ r(\vec{p})<\infty\right\},\quad{Q_{\mathrm{ref}}=\left\{(\vec{p},r(\vec{p}):\vec{p}\in P_{\circ},\ r(\vec{p})>1\right\}}.$
@@
$\left\|T\right\|_{\infty,\mathrm{BMO}}=\sup_{(f_{1},\ldots,f_{m})\in\mathcal{B}^{\vec{p}}}\left\|T(f_{1},\ldots,f_{m})\right\|_{\mathrm{BMO}(\mathbb{R}^{d})},\quad\vec{p}=(\infty,\ldots,\infty),$
where the $\mathrm{BMO}(\mathbb{R}^{d})$ norm is defined by
(2.8) $\left\|f\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}=\sup_{Q\,\mathrm{cube}}\frac{1}{|Q|}\int_{Q}\left|f(x)-f_{Q}\right|\,{\rm d}x,\quad f_{Q}=\frac{1}{|Q|}\int_{Q}f(y)\,{\rm d}y.$
@@
For each $(\vec{p},r)\in Q$, there exists $C_{\vec{p},r}>0$ such that for any $T$ such that $U(f_{0},\ldots,f_{m})=\left\langle f_{0},T(f_{1},\ldots,f_{m})\right\rangle$ is a bounded wavelet form,
@@
Furthermore, there exists $C_{\infty}>0$ and $\bar{T}$ an extension of $T$ such that
(2.10) $\left\|\bar{T}\right\|_{\infty,\mathrm{BMO}}\leq C_{\infty}\sup_{z\in Z^{d}}\left|\varepsilon_{z}\right|.$
@@ 3. Wavelet representation of compact Calderón-Zygmund forms
(3.1) $|K(x_{0},\ldots,x_{m})|\leq\dfrac{C_{K}}{\left|x_{0}-\mathbf{x}\right|_{2}^{md}},\quad\mathbf{x}=(x_{1},\ldots,x_{m}),$
(3.2) $\max_{j=0,\ldots,m}|\Delta^{j}_{h}K(x_{0},\ldots,x_{m})|\leq\dfrac{C_{K}|h|^{\delta}}{\left|x_{0}-\mathbf{x}\right|_{2}^{md+\delta}},$
where $\Delta_{h}^{j}$ denotes the difference operator in the $j$-th position , and $\left|\cdot\right|_{2}$ is defined in (2.1). Let $\Lambda:\times_{j=0}^{m}{\mathcal{S}}(\mathbb{R}^{d})\to\mathbb{C}$ be a continuous $(m+1)$-linear form. We say $\Lambda$ is an $(m+1)$-linear singular integral form if there exists a singular integral kernel $K$ such that for any $f_{0},\ldots,f_{m}\in{\mathcal{S}}(\mathbb{R}^{d})$ with $\cap_{j=0}^{m}{\mathrm{supp}\,}f_{j}=\varnothing$ one has
@@
Let ${\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ be the space of all continuous functions $f$ on $\mathbb{R}^{d}$ for which $\lim_{|x|\to\infty}f(x)=0$. Then, the Banach space of functions with continuous mean oscillation, $\mathrm{CMO}(\mathbb{R}^{d})$, is defined to be the closure of ${\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ in the norm $\left\|\cdot\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}$ , defined in (2.8) above.
@@
and furthermore, there exist $\mathfrak{b}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$ such that (1.2) holds. The rigorous definition of (1.2) is as follows. Let $\theta\in{\mathrm{C}}^{\infty}_{0}(\mathbb{R}^{d})$ with $\theta(x)=1$ for $|x|\leq 1$. Then, for $t>0$, set $\theta_{t}=\theta(\cdot t)$. We say (1.2) holds if for all $\varphi\in{\mathcal{S}}^{0}(\mathbb{R}^{d})$,
@@
If $T$ is a continuous $m$-linear operator from ×$\times_{j=1}^{m}{\mathcal{S}}\to{\mathcal{S}}^{\prime}$ such that $\left\langle f_{0},T(f_{1},\ldots,f_{m})\right\rangle=\Lambda(f_{0},\ldots,f_{m})$ for some CZ form $\Lambda$, then $T$ is said to be a Calderón-Zygmund operator (CZO). Furthermore, let $\phi$ be a mother wavelet and $\psi^{2},\psi^{4}$ be the Schwartz functions from Lemma 2.4. For $j=1,\ldots,m$ define